66 randomly selected students were asked the number of pairs of shoes they have. Let X represent the number of pairs of shoes. The results are as follows:

# of Pairs of Shoes 4 5 6 7 8 9 10 11 12 13 14 15
Frequency 5 5 3 5 5 6 7 5 7 5 6 7



Round all your answers to 4 decimal places where possible.

The mean is: _________

The median is: _________

The sample standard deviation is: ________  

The first quartile is: _______

The third quartile is: _______

What percent of the respondents have at least 13 pairs of Shoes? ______%

16% of all respondents have fewer than how many pairs of Shoes? ________
Answer #1

The mean is =9.9242

The median is =npth value =66*0.5 th value =33th value =10

sample standard deviation =3.4477

first quartile is 7

third quartile is 13

percent of the respondents have at least 13 pairs of Shoes =(18/66)*100 =27.2727%

16% of all respondents have fewer than how many pairs of Shoes =5